# Shortest Paths in a Weighted Simplicial Complex

### v1

@@ 3 Mark all vertices as unvisited
5while there exists an unvisited $(d-1)$-simplex $\sigma_{1}$ with $t(\sigma_{1})<\infty$ do
@@ 10      Mark $\sigma_{1}$ as visited
11
12return Distance function $t\colon X_{(d-1)}\to[0,\infty]\,$, where $t(\sigma)$ is the minimum total weight of a chain of $d$-simplices from $\sigma_{0}$ to $\sigma$.
13
@@ 3.1 Analysis of Algorithm 1
Given a weighted $d$-complex with nonnegative weight function $w$ and a starting $(d-1)$-simplex $\sigma_{0}$, Algorithm 1 computes $d_{w}(\sigma_{0},\sigma_{1})$ for every $\sigma_{1}\in X_{(d-1)}.$
Proof
@@
We prove that Algorithm 1 correctly computes the shortest $d$-path from the source $(d-1)$-simplex $\sigma_{0}$ to every other $(d-1)$-simplex in a weighted complex $(X,w)$, where $w$ is the weight function on the edges. We use Mathematical Induction to show that at each step of Dijkstra’s algorithm, the distance $t(\sigma_{1})$ is always the correct shortest $d$-path from the source $\sigma_{0}$ to $(d-1)$-simplex $\sigma_{1}$ for all vertices $\sigma_{1}\in X_{(d-1)}$.
@@
At any step of the algorithm, let $\mathcal{S}$ be the set of $(d-1)$-simplexes whose shortest distance from the source $\sigma_{0}$ has been finalised. Initially, the set $\mathcal{S}=\{\sigma_{0}\}$, and the algorithm sets $t(\sigma_{0})=0$ and $t(\sigma_{2})=\infty$ for all $\sigma_{2}\neq\sigma_{0}$. Clearly, the distance from $\sigma_{0}$ to $\sigma_{0}$ is correct since $t(\sigma_{0})=0$. So, the algorithm correctly initialises the distances.
@@
Assume that the algorithm has correctly computed the shortest $d$-path distances for all vertices in $\mathcal{S}$, where $|\mathcal{S}|=m$ for some $m\in\mathbb{N}$. Let $\sigma_{1}$ be the next $(d-1)$-simplex added to $\mathcal{S}$. By the algorithm’s rule, $\sigma_{1}$ is the unvisited $(d-1)$-simplex with the smallest tentative distance $t(\sigma_{1})$. We need to show that this choice of $\sigma_{1}$ guarantees that $t(\sigma_{1})$ is the correct shortest $d$-path from $\sigma_{0}$ to $\sigma_{1}$.
@@
Any $d$-path $P$ from $\sigma_{0}$ to $\sigma_{1}$ must exit $\mathcal{S}$ before reaching $\sigma_{1}$. The induction hypothesis states that the length of the shortest $d$-path from $S$ to $\sigma_{1}$ is $t(\sigma_{1})$. The induction hypothesis and the choice of $\sigma_{1}$ also guarantee that a $d$-path visiting any $(d-1)$-simplex outside $\mathcal{S}$ and later reaching $\sigma_{1}$ has length at least $t(\sigma_{1})$. Hence $d_{w}(\sigma_{0},\sigma_{1})=t(\sigma_{1})$. Therefore, the algorithm correctly computes the shortest $d$-path to $\sigma_{1}$ and adds $\sigma_{1}$ to $\mathcal{S}$. The algorithm then proceeds to the next $(d-1)$-simplex in the same way, ensuring that at every step, the shortest $d$-paths from $\sigma_{0}$ to all vertices in $\mathcal{S}$ are correctly computed.
@@
The algorithm terminates when $\mathcal{S}=X_{d-1}$, at which point $t(\sigma_{1})$ for each $(d-1)$-simplex $\sigma_{1}$ in $X_{d-1}$ represents the shortest $d$-path from $\sigma_{0}$ to $\sigma_{1}$.
@@
Since the algorithm always selects the $(d-1)$-simplex with the smallest tentative distance at each step and updates distances optimally, the final distances $t(\sigma_{1})$ are indeed the correct shortest $d$-path distances from $\sigma_{0}$ to every $(d-1)$-simplex in $X_{d-1}$.
Hence, the correctness of the algorithm Algorithm 1 follows from Theorem 3.1.
Complexity.
@@
Now we deduce the time and memory complexity of Algorithm 1. For clarity we denote the number of $d-1$-simplices by $n=|X_{d-1}|$ and the number of $d$-simplices by $m=|X_{d}|$. We will analyse the algorithm assuming that we use the binary heap data structure to store the distances. Now, if we use a binary heap structure for storing the current distances, then scanning the binary heap to find the smallest distance (extract-min) would take $\mathcal{O}(\log n)$ operations per extraction. Next, updating the entry (decrease-key) would take $\mathcal{O}(\log n)$ operations per iteration. Hence, the time-complexity of the algorithm is a sum of $n$ extractions and up to $m$ decrease-key operations. Therefore, the overall runtime of the algorithm is given by $\mathcal{O}((n+m)\log n)$.
@@
For each $d-1$-simplex, we store the tentative distance from the starting point, which requires $\mathcal{O}(n)$ memory. Next, for each $d-1$-simplex, we need to store a boolean value that represents whether that simplex has been visited or not. This again requires $\mathcal{O}(n)$ memory. The algorithm also uses a priority queue to quickly extract an unvisited $d-1$-simplex with the smallest distance. The queue stores up to $n$ entries at once, each corresponding to a $d-1$-simplex and its associated distance. Therefore, the queue requires $\mathcal{O}(n)$ memory. Now, to find the neighbours of a given $d-1$-simplex, we must know all the $d$-simplices that contain the given $d-1$-simplex. It requires $\mathcal{O}(m)$ memory to store this adjacency information. This makes the total memory requirement of the algorithm to be $\mathcal{O}(m+n)$.
@@ 4 Conclusion
In this study, we have defined the concept of weight in a $d$-complex to facilitate our study of $d$-complex as an object in discrete mathematics and graph theory rather than a topological object. The concept of a $d$-path comes from this naturally. The main contribution of our paper is our novel algorithm to find the shortest path in a weighted $d$-complex. With its vast application in various fields, our algorithm is poised to play a crucial role in the study of complexes as mathematical objects with real-life applications. Our algorithm can be seen as the $d$-complex version of Dijkstra’s Algorithm for graphs. Analogy of a weighted $d$-complex with a weighted graph opens numerous questions in this field of study. The concepts introduced in this paper pave the path for a line of study in which simplicial complexes will be viewed as geometric computational objects rather than abstract entities.
This work extends beyond just providing a thorough analysis of our algorithm, this also indicates possible integration of simplicial complexes with various concepts of graph theory, specially weighted graphs.
One future aspect of study is the concept of a $d$-cycle (see Definition 3). The natural question to ask is, given $d$-complex how can we decide if the given $d$-complex contains a $d$-cycle or not. Now, if a $d$-cycle exists, how does one identify a $d$-cycle? The mono-dimensional version of these questions are of immense interest in the study of theoretical computer science and graph theory. The questions that we raised here are beyond the scope of this paper, hence we leave them as future directions.

### v2

@@ 3 Mark all vertices as unvisited
5while there exists an unvisited (d1)(d-1)-simplex $\sigma_{1}$ with $t(\sigma_{1})<\infty$ do
@@ 11  Mark $\sigma_{1}$ as visited
12
13 end while
14return Distance function $t\colon X_{(d-1)}\to[0,\infty]\,$, where $t(\sigma)$ is the minimum total weight of a chain of $d$-simplices from $\sigma_{0}$ to $\sigma$.
15
@@ 4.1 Analysis of Algorithm 3
Given a weighted $d$-complex with non-negative weight function $w$ and a starting (d1)(d-1)-simplex $\sigma_{0}$, Algorithm 3 computes $d_{w}(\sigma_{0},\sigma_{1})$ for every $\sigma_{1}\in X_{(d-1)}.$
Proof.
@@
We prove that Algorithm 3 correctly computes the shortest $d$-path from the source (d1)(d-1)-simplex $\sigma_{0}$ to every other (d1)(d-1)-simplex in a weighted complex $(X,w)$, where $w$ is the weight function on the edges. We use Mathematical Induction to show that at each step of Dijkstra’s algorithm, the distance $t(\sigma_{1})$ is always the correct shortest $d$-path from the source $\sigma_{0}$ to (d1)(d-1)-simplex $\sigma_{1}$ for all vertices $\sigma_{1}\in X_{(d-1)}$.
@@
At any step of the algorithm, let $\mathcal{S}$ be the set of (d1)(d-1)-simplexes whose shortest distance from the source $\sigma_{0}$ has been finalised. Initially, the set $\mathcal{S}=\{\sigma_{0}\}$, and the algorithm sets $t(\sigma_{0})=0$ and $t(\sigma_{2})=\infty$ for all $\sigma_{2}\neq\sigma_{0}$. Clearly, the distance from $\sigma_{0}$ to $\sigma_{0}$ is correct since $t(\sigma_{0})=0$. So, the algorithm correctly initialises the distances.
@@
Assume that the algorithm has correctly computed the shortest $d$-path distances for all vertices in $\mathcal{S}$, where $|\mathcal{S}|=m$ for some $m\in\mathbb{N}$. Let $\sigma_{1}$ be the next (d1)(d-1)-simplex added to $\mathcal{S}$. By the algorithm’s rule, $\sigma_{1}$ is the unvisited (d1)(d-1)-simplex with the smallest tentative distance $t(\sigma_{1})$. We need to show that this choice of $\sigma_{1}$ guarantees that $t(\sigma_{1})$ is the correct shortest $d$-path from $\sigma_{0}$ to $\sigma_{1}$.
@@
Any $d$-path $P$ from $\sigma_{0}$ to $\sigma_{1}$ must exit $\mathcal{S}$ before reaching $\sigma_{1}$. The induction hypothesis states that the length of the shortest $d$-path from $S$ to $\sigma_{1}$ is $t(\sigma_{1})$. The induction hypothesis and the choice of $\sigma_{1}$ also guarantee that a $d$-path visiting any (d1)(d-1)-simplex outside $\mathcal{S}$ and later reaching $\sigma_{1}$ has length at least $t(\sigma_{1})$. Hence $d_{w}(\sigma_{0},\sigma_{1})=t(\sigma_{1})$. Therefore, the algorithm correctly computes the shortest $d$-path to $\sigma_{1}$ and adds $\sigma_{1}$ to $\mathcal{S}$. The algorithm then proceeds to the next (d1)(d-1)-simplex in the same way, ensuring that at every step, the shortest $d$-paths from $\sigma_{0}$ to all vertices in $\mathcal{S}$ are correctly computed.
@@
The algorithm terminates when $\mathcal{S}=X_{d-1}$, at which point $t(\sigma_{1})$ for each (d1)(d-1)-simplex $\sigma_{1}$ in $X_{d-1}$ represents the shortest $d$-path from $\sigma_{0}$ to $\sigma_{1}$.
@@
Since the algorithm always selects the (d1)(d-1)-simplex with the smallest tentative distance at each step and updates distances optimally, the final distances $t(\sigma_{1})$ are indeed the correct shortest $d$-path distances from $\sigma_{0}$ to every (d1)(d-1)-simplex in $X_{d-1}$. ∎
Hence, the correctness of the algorithm Algorithm 3 follows from Theorem 1.
Complexity
@@
Now we deduce the time and memory complexity of Algorithm 3. For clarity we denote the number of d1d-1-simplices by $n=|X_{d-1}|$ and the number of $d$-simplices by $m=|X_{d}|$. We will analyse the algorithm assuming that we use the binary heap data structure to store the distances. Now, if we use a binary heap structure for storing the current distances, then scanning the binary heap to find the smallest distance (extract-min) would take $\mathcal{O}(\log n)$ operations per extraction. Next, updating the entry (decrease-key) would take $\mathcal{O}(\log n)$ operations per iteration. Hence, the time-complexity of the algorithm is a sum of $n$ extractions and up to $m$ decrease-key operations. Therefore, the overall runtime of the algorithm is given by $\mathcal{O}((n+m)\log n)$.
@@
For each d1d-1-simplex, we store the tentative distance from the starting point, which requires $\mathcal{O}(n)$ memory. Next, for each d1d-1-simplex, we need to store a boolean value that represents whether that simplex has been visited or not. This again requires $\mathcal{O}(n)$ memory. The algorithm also uses a priority queue to quickly extract an unvisited d1d-1-simplex with the smallest distance. The queue stores up to $n$ entries at once, each corresponding to a d1d-1-simplex and its associated distance. Therefore, the queue requires $\mathcal{O}(n)$ memory. Now, to find the neighbours of a given d1d-1-simplex, we must know all the $d$-simplices that contain the given d1d-1-simplex. It requires $\mathcal{O}(m)$ memory to store this adjacency information. This makes the total memory requirement of the algorithm to be $\mathcal{O}(m+n)$.
5 Generalization of the Bellman-Ford Algorithm
We now present a relaxation-based algorithm that generalizes the classical Bellman-Ford procedure to the higher-dimensional setup introduced in this paper. In contrast to Algorithm 3, which assumes non-negative $d$-simplex weights, the following method correctly computes shortest weighted (d1)(d-1)-paths even in the presence of negative weights and detects the existence of negative $d$-cycles.
Let $(X,w)$ be a weighted $d$-complex with the vertex set $V:=X_{d-1}$ and the adjacency relation defined as before: two (d1)(d-1)-simplices $\sigma,\sigma^{\prime}\in V$ are neighbors if and only if $\sigma\cup\sigma^{\prime}\in X_{d}$. The cost associated with the pair $(\sigma,\sigma^{\prime})$ is the weight $w(\sigma\cup\sigma^{\prime})$ of that $d$-simplex. Define the edge set
$E:=\{(\sigma,\sigma^{\prime})\in V\times V\mid\sigma\cup\sigma^{\prime}\in X_{d}\},\qquad c(\sigma,\sigma^{\prime}):=w(\sigma\cup\sigma^{\prime}).$
A walk of (d1)(d-1)-simplices $\Pi=(\sigma_{0},\sigma_{1},\dots,\sigma_{\ell})$ with $(\sigma_{i-1},\sigma_{i})\in E$ for all $i$ is called a $d$-path, and its total weight is $w(\Pi):=\sum_{i=1}^{\ell}c(\sigma_{i-1},\sigma_{i}).$ A negative $d$-cycle is a closed $d$-path $\Pi$ with $w(\Pi)<0$.
Input : Weighted $d$-complex $(X,w)$, source $\sigma_{0}\in X_{d-1}$
Output : Shortest distances $t(\sigma)$ from $\sigma_{0}$; or report a negative $d$-cycle
1
[…]
3
4foreach $\sigma\in V$ do
5 $t(\sigma)\leftarrow+\infty$
6 end foreach
[…]
8
9for $i\leftarrow 1$ to |V|1|V|-1 do
10 foreach $(\sigma,\sigma^{\prime})\in E$ do
11    if $t(\sigma)\neq+\infty$ and $t(\sigma^{\prime})>t(\sigma)+c(\sigma,\sigma^{\prime})$ then
12       $t(\sigma^{\prime})\leftarrow t(\sigma)+c(\sigma,\sigma^{\prime})$
13
14      end if
15
16   end foreach
17
18 end for
19
[…]
21 foreach $(\sigma,\sigma^{\prime})\in E$ do
22 if $t(\sigma)\neq+\infty$ and $t(\sigma^{\prime})>t(\sigma)+c(\sigma,\sigma^{\prime})$ then
23    $U\leftarrow U\cup\{\sigma^{\prime}\}$
24
25   end if
26
27 end foreach
28
291exif $U\neq\varnothing$ then
30 Report: “Negative $d$-cycle detected” and return the set of all (d1)(d-1)-simplices reachable from $U$ under $E$
31
32 end if
33else
34 return $t(\cdot)$
35
36 end if
Algorithm 4 Bellman-Ford on Weighted $d$-Complexes (BF-$d$)
Discussion
Algorithm 4 preserves the same combinatorial interpretation as Algorithm 3: relaxation is performed over the adjacency induced by $X_{d}$, but instead of greedily fixing vertices, all edges are relaxed iteratively. The procedure tolerates negative simplex weights and identifies negative $d$-cycles via the standard one-pass verification step.
Example 1 (Negative $d$-Cycle Detection).
Consider a $2$-complex $X$ whose $1$-simplices (edges) $\{a,b,c\}$ bound two $2$-simplices $\tau_{1},\tau_{2}$ sharing a common edge $b$. Assign weights $w(\tau_{1})=-1$, $w(\tau_{2})=-2$. The adjacency relation on $V=X_{1}=\{a,b,c\}$ contains the edges $(a,b)$, $(b,c)$, $(c,a)$ with total cycle weight $w(\tau_{1})+w(\tau_{2})=-3<0$. Executing Algorithm 4 with source $a$ yields decreasing distances after |V|1|V|-1 relaxations, and the final verification pass identifies $a,b,c\in U$, signalling a negative $2$-cycle corresponding to the two faces $\tau_{1},\tau_{2}$.
Lemma 5 (Correctness of BF-$d$).
Let $(X,w)$ be a finite weighted $d$-complex and let $\sigma_{0}\in X_{d-1}$. If $(X,w)$ contains no negative $d$-cycle, then upon termination Algorithm 4 satisfies
$t(\sigma)=\min_{\Pi:\,\sigma_{0}\leadsto\sigma}w(\Pi)\quad\text{for all }\sigma\in X_{d-1},$
where the minimum is taken over all $d$-paths from $\sigma_{0}$ to $\sigma$. If $(X,w)$ contains a negative $d$-cycle, then $U\neq\varnothing$ after the final iteration, and every element of $U$ lies on or is reachable from some negative $d$-cycle.
Proof.
The argument is identical to the classical Bellman-Ford invariant proof, with the graph of (d1)(d-1)-simplices replacing the vertex set. Let $\Pi_{k}(\sigma)$ denote the family of $d$-paths from $\sigma_{0}$ to $\sigma$ containing at most $k$ edges. We prove by induction on $k$ that after the $k$-th relaxation round, $t(\sigma)\leq\min_{\Pi\in\Pi_{k}(\sigma)}w(\Pi)$ for all $\sigma$. For $k=0$ this is immediate. For the inductive step, relaxing each $(\sigma^{\prime},\sigma)\in E$ allows one additional edge, so $t(\sigma)\leq t(\sigma^{\prime})+c(\sigma^{\prime},\sigma)\leq\min_{\Pi\in\Pi_{k-1}(\sigma^{\prime})}w(\Pi)+c(\sigma^{\prime},\sigma)=\min_{\Pi\in\Pi_{k}(\sigma)}w(\Pi)$. After |V|1|V|-1 rounds, every simple path (of length at most |V|1|V|-1) has been considered. If no negative cycle exists, a shortest simple path attains the optimum. If a negative cycle exists, an additional relaxation decreases $t(\sigma)$ for some $(\sigma,\sigma^{\prime})$, placing $\sigma^{\prime}\in U$. Thus the lemma follows. ∎
Complexity
Let $n:=|X_{d-1}|$ and $m_{d}:=|X_{d}|$. Each $d$-simplex induces at most $(d+1)d$ neighbor pairs. Denoting $m_{e}:=|E|=\Theta(d^{2}m_{d})$, Algorithm 4 performs $\mathcal{O}(n)$ passes over $m_{e}$ edges, giving a total time complexity $\mathcal{O}(n\cdot m_{e})=\mathcal{O}(n\,d^{2}m_{d})$ and space complexity $\mathcal{O}(n+m_{d})$.
6 Real-life Example of Our Setup
In this section we provide a real-life example of our setup that shows how our setup can better model some of the real-life problems than graphs.
In modern communication systems, higher-order interactions often arise naturally, particularly in multi-agent or wireless sensor networks, where groups of devices collaborate to sense, process, or transmit data. These scenarios go beyond the pairwise connectivity that ordinary graphs can express, and are therefore more accurately represented by simplicial complexes.
Consider a distributed network of sensor nodes or autonomous drones deployed over a geographical region. Each sensor node represents a 0-simplex. Two nodes that can communicate directly via a wireless link form a $1$-simplex. Moreover, subsets of three or more nodes that are capable of joint sensing or cooperative transmission form higher-dimensional simplices. Thus, a 2-simplex $\{i,j,k\}$ indicates a cooperative triplet of sensors that can jointly process or relay data, and more generally, a $d$-simplex represents a $d$-agent coalition capable of group-level communication or computation.
In this setting, a weighted $d$-complex $(X,w)$ arises naturally, where each $d$-simplex $\tau\in X_{d}$ carries a weight $w(\tau)$ representing the total energy cost, delay, or bandwidth consumption associated with that cooperative operation. Two (d1)(d-1)-simplices $\sigma,\sigma^{\prime}$ are said to be adjacent if $\sigma\cup\sigma^{\prime}\in X_{d}$, i.e., if they share participation in some $d$-agent transmission. The communication process from one coalition of agents to another can therefore be viewed as a $d$-path in the complex.
The algorithms introduced in this paper can then be interpreted in this framework. The generalized BFS and DFS algorithms on $d$-complexes discover all clusters of agents reachable under cooperative adjacency, thereby serving as topological discovery tools for multi-agent connectivity. The simplicial Dijkstra algorithm (see Algorithm 3) computes the minimum-cost cooperative route for data transmission from one agent group to another, optimizing total communication energy or latency across multiple overlapping coalitions. The generalized Bellman–Ford procedure (see Algorithm 4) extends this to time-varying or uncertain environments where communication costs may fluctuate or even become negative (e.g., due to energy harvesting or incentive mechanisms), while also detecting unstable feedback loops corresponding to negative $d$-cycles.
In summary, the shortest $d$-path problem in weighted simplicial complexes provides a rigorous combinatorial model for cooperative routing in multi-agent systems. By extending classical shortest-path computations to higher-dimensional adjacency, this framework captures collective communication patterns that cannot be modeled by graphs alone. Consequently, the proposed algorithms have potential applications in wireless sensor networks, distributed robotics, and cooperative multi-agent systems where higher-order interactions play a central role.
@@ 7 Conclusion
In this study, we have defined the concept of weight in a $d$-complex to facilitate our study of $d$-complex as an object in discrete mathematics and graph theory rather than a topological object. The concept of a $d$-path comes from this naturally. The main contributions of our paper are our novel algorithms to find the shortest path in a weighted $d$-complex. With its vast application in various fields, our algorithms are poised to play a crucial role in the study of complexes as mathematical objects with real-life applications. Our algorithms can be seen as the $d$-complex version of Dijkstra’s Algorithm and Bellman-Ford Algorithm for graphs. Analogy of a weighted $d$-complex with a weighted graph opens numerous questions in this field of study. The concepts introduced in this paper pave the path for a line of study in which simplicial complexes will be viewed as geometric computational objects rather than abstract entities.
This work extends beyond just providing a thorough analysis of our algorithms, this also indicates possible integration of simplicial complexes with various concepts of graph theory, specially weighted graphs.
One future aspect of study is the concept of a spanning $d$-tree, where similar to graphs, the case of absence of a cycle in a connected simplicial complex can be defined to be a $d$-tree. The natural question to ask is, given $d$-complex how can we decide if the given $d$-complex is $d$-tree or not. Furthermore, if a $d$-complex is given, how does one identify a minimal spanning $d$-tree? The mono-dimensional version of these questions are of immense interest in the study of theoretical computer science and graph theory, which make these questions raised above worth pursuing.
8 Declarations
Funding
No funding was received for conducting this study.
Conflict of Interest
The authors declare that they have no potential conflict of interest.
Data Availability
No datasets were generated or analyzed during the current study.